# Dynamic tariffs-based demand response in retail electricity market under uncertainty

### v5

@@ 3 Problem formulation
Most electricity markets exhibit a hierarchical decision-making structure, that allows participants with market power to potentially anticipate the reactions of subsequent decision-making entities. In this regard, bilevel programming is an adequate optimization-based modeling technique to represent this sort of interaction. It can be defined in various forms, including single-leader single-follower, single-leader multi-follower, multi-leader single-follower, and multi-leader multi-follower games. While the single leader-single follower case has been extensively analyzed since von Stackelberg’s seminal work (Von Stackelberg, 1952), its extensions, accounting for multiple followers, remain less explored. Additionally, there aren’t general numerical methods for finding equilibrium solutions efficiently, and the existing algorithms are problem-specific.
@@
[…]
@@
$\lambda_{j\omega}$ total quantity variation of demand for consumer $j$ at scenario $\omega$
@@
In this work, we consider a single retailer who wants to optimize its expected profit, and $\left|J\right|$ consumers who want to minimize the difference between the cost of purchasing electricity and the utility brought by their consumption. That is, the retailer determines the selling prices to maximize its expected profit, which depends on the consumers’ purchases and the consumers determine their load demands, which depend on the retail price tariffs. The consumers want to minimize the disutility costs. First, we develop a bi-level programming problem assuming the retailers behave as leaders (upper-level) and the consumers as followers (lower-level). Then, for comparison purposes, we analyze a competitive equilibrium between the retailer and consumers, assuming no market power.
@@
The bi-level programming with a strategic retailer, is reformulated as an MPEC problem. This is done by recasting the consumers’ problems using their KKT optimality conditions, which are introduced as constraints in the retailer’s problem. Uncertainties related to consumers’ demand, spot market quantities, and prices are addressed using stochastic programming.
Even though electricity generation costs are approximated by quadratic production costs, consumer utilities are frequently simplified with linear functions. Quadratic utility functions are interesting in DR programs because of their non-decreasing marginal benefits (Niromandfam et al., 2020; Chen et al., 2010). To demonstrate that the overall implications of the proposed models are not solely dependent on a specific form of the utility function, we employ both the linear utility function and its quadratic extension, which is also a popular form for describing customer behavior using elasticity (Niromandfam et al., 2020).
Due to the assumption that the model provides no possibility for consumers to switch to a different retailer (short-term model), the utility function prevents the retailer from increasing the price tariffs possibly up to infinity to maximize its profit [^1] To handle such challenges, for example, Zugno et al. (2013) include minimum, maximum, and average prices into the consumers’ problem as constraints..
@@
For the equilibrium model with perfect competition, we formulate a MILP problem by linearizing the complementarity conditions of the corresponding retailer and consumers’ KKT optimality conditions. In addition, an NLP reformulation of the equilibrium problem is derived using the methods proposed by Leyffer and Munson (2010), which provides a solution equivalent to the MILP formulation.
The model has first-stage and second-stage decisions. Prior to the day-ahead market clearance, the retailer decides the price tariff for hour $t$ (first stage). However, the spot market price, the quantity purchased by the retailer from the electricity market, the quantity bought by consumers, and the electricity quantity imbalance decisions are second-stage and scenario-dependent variables.
@@ 3.1 Retailer’s problem
The retailer maximizes her expected profit subject to technical and economic constraints, and accounting for uncertainty, which is modeled by a discrete number of scenarios. The retailer’s problem is stated mathematically as follows:
@@
where $\bf{X}$=$\{P_{t},q_{t\omega}^{S},\delta_{t\omega},y_{t\omega}\}$ is the retailer’s set of decision variables, while $\mu_{t\omega},\alpha_{t\omega}^{+},\alpha_{t\omega}^{-}$, $\theta_{t\omega},\gamma_{t},\beta_{t\omega}$ are the dual variables associated with the constraints ((3.1b))-((3.1g)).
The first term in the objective function is the revenue for the retailer from the quantity bought by consumer $j$, at hour $t$, with scenario $\omega$ and price $P_{t}$, which is the retail price of electricity (tariff), a first-stage decision determined by the retailer for each hour $t$. The second term is the cost of purchasing the electricity quantity $q_{t\omega}^{S}$ at the spot market price $P_{t\omega}^{S}$ at hour $t$ and with scenario $\omega$. The last term is a cost (penalization of electricity imbalances), which can be positive or negative (variable $y_{t\omega}$ represents the absolute value of deviations of the electricity imbalances), and $\sigma_{\omega}$ is the probability associated with scenario $\omega$. The summation of the product over index $\omega\in\Omega$ represents the expected profit, which is what the retailer wants to maximize over the time horizon.
Constraint ((3.1b)) is the electricity balancing constraint that guarantees the total demand by all consumers at hour $t$ with scenario $\omega$ minus retailer’s supply is the imbalance at that particular time. The linear set of constraints ((3.1c)), ((3.1d)) and ((3.1e)) ensure that $y_{t\omega}=|\delta_{t\omega}|$. These imbalances are penalized in the objective function with weight $C$. Constraints ((3.1f)) and ((3.1g)) are non-negativity constraints on the spot market quantity and price decided by the retailer for tariff at hour $t$, respectively. $P_{t}$ is a first-stage decision, which needs to be settled before the uncertainty realization, while $q_{jt\omega},q_{t\omega}^{S},\delta_{t\omega}$, $P_{t\omega}^{S}$ and $y_{t\omega}$ are second stage decisions. Correspondingly, the dual variables for each group of constraints are indicated on the right side of a colon.
@@ 3.2 Consumers’ problem
In modeling the consumers’ problem with DR, it is common to express the type of consumers (residential, industrial, or commercial) explicitly so that part of their consumption is considered flexible, and part of the consumption is inelastic (demand that must be met strictly at all times). For instance, Halvgaard et al. (2012) model heat pumps for heating residential buildings, where the heating system of the house becomes a flexible consumption in the smart grid. Zugno et al. (2013) extend the above model by considering the heating dynamics of a building, indoor temperature, or temperature inside a water tank where flexibility (the output of interest) is characterized through the indoor temperature. They also consider the inflexible part of consumption in their model. However, in our model setting, including individual appliances would require a large number of new variables and constraints, which will further complicate its numerical resolution and the economical interpretation of the results that we aspire to deliver. Moreover, we believe that the main demand patterns/profiles can be already accounted for by the appropriate inter-temporal adjustment of the utility function parameters.
By considering the consumers’ utility function explicitly in the objective function, we explore a residential consumers’ model with flexibility towards the dynamic price reported by the retailer. Thus, the consumers face an economic problem where it is crucial to balance the trade-off between the cost of electricity procurement and the discomfort of deviating from the utility of consumption given a flexible price and the possibility to shift the consumption to low price periods. In particular, the consumers’ problem can be represented within a game-theoretical setting as it is parametrized by the decision of the retailer (hourly price $P_{t}$). The problem for consumer $j$ with scenario $\omega$ can be expressed as:
@@
where the objective function is the minus social welfare (cost minus utility) and $q_{jt\omega}$ (inelastic demand) and $\Delta_{jt\omega}^{C}$ (flexible demand) are consumer’s decision variables. In the objective function $P_{t}q_{jt\omega}$ is the cost of purchasing $q_{jt\omega}$ electricity and $A_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})-\frac{1}{2}B_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})^{2}$ is the quadratic utility function that measures the benefit consumer $j$ achieves by consuming the amount of energy ($q_{jt\omega}+\Delta_{jt\omega}^{C}$) during hour $t$, scenario $\omega$.
Considering the consumers’ objective function without any further constraints, consumption would only take place in periods when the real-time price is lower than the marginal benefit. Note that the marginal utility is a demand function $P_{t}=A_{jt\omega}-B_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})$ where electricity quantity demanded decreases as electricity price increases. In other words, at the optimum the marginal utility (benefit) from consuming electricity equals the retail price tariff $P_{t}$.
Parameters $A_{jt\omega}$ and $B_{jt\omega}$ are the intercept and slope of the demand function, which are important parameters to characterize the behavior of consumers toward the market outcomes of the model. Different values for these parameters can capture the dynamics of consumers’ demand. Constraint ((3.2c)) sets the lower and upper limits for flexible consumption, where $\Delta_{j}^{C_{max}}$ is the level of flexibility of consumer $j$. The inequalities guarantee that the flexible part of consumption at hour $t$ falls in the range between $-\Delta_{j}^{C_{max}}$ and $+\Delta_{j}^{C_{max}}$, which bound consumption increase and decrease, respectively. If $\Delta_{j}^{C_{max}}=0$, there is no flexibility in demand, while $\Delta_{j}^{C_{max}}>0$, corresponds to the flexible demand. In the latter case, the consumer considers a dynamic price tariff that ensures cost savings and better welfare. Besides, constraint ((3.2d)) ensures the net total flexible demand during the planning time horizon is zero. In the consumers’ problem, $q_{jt\omega}$ and $\Delta_{jt\omega}^{C}$ are second-stage decisions. Note that since the dynamic electricity price $P_{t}$ enters the consumers’ problem as a parameter (it is only a variable in the retailer problem), the optimization problems of the consumers are convex (quadratic objective function with linear constraints). Therefore, we can replace each consumer’s problem with its corresponding KKT optimality conditions (Haghighat and Kennedy, 2012), which are sufficient for optimality.
@@ 3.2.1 KKT formulation of the consumer problem
where $\mathcal{L}_{j\omega}$ stands for the Lagrangian function for consumer $j$ and scenario $\omega$ of problem ((3.2)). By convention, the symbol $\perp$ indicates complementarity so that any of the two inequalities is satisfied as equality, i.e., the product of each expression and the corresponding dual variable must be zero. Note that the system of KKT optimality conditions is linear, with the exception of the complementarity conditions ((3.3d))-((3.3f)).
@@ 3.3 The Retailer’s MPEC problem
Figure 2: Game-theoretical framework for modeling demand response as a single leader-followers problem. Solid arrows indicate information signals from the retailer, while dashed arrows represent inferences on the consumers’ behavior.
@@
The second type of nonlinearities can be linearized using the strategy proposed in Fortuny-Amat and McCarl (1981), which is by applying the big-M technique. However, we still face the nonlinear product $P_{t}q_{jt\omega}$ which, as opposed to other problem settings Ruiz and Conejo (2009) and Sadat and Fan (2017) cannot be linearized by employing the strong duality theorem. However, due to the observed good numerical performance, we have opted to solve the original MPEC problem with a state-of-the-art NLP solver like KNITRO (Artelys KNITRO, 2021).
@@ 3.4 The equilibrium model
Considering the perfect competition, we tackle the equilibrium model by jointly solving retailers’ and consumers’ KKT optimality conditions using two techniques: i) by linearizing and casting the KKT systems of equations as a MILP problem employing the Fortuny-Amat and McCarl (1981) big-M technique, and ii) by directly addressing and reformulating the problem as a NLP using the approach in Leyffer and Munson (2010). However, we mainly present the reformulation of the equilibrium problem as a MILP and its results, for our analysis. For details on the reformulation of NLP problems, the interested reader is referred to Abate et al. (2021).
@@ 3.4.1 KKT formulation of the retailer problem
To derive the retailer’s KKT optimality conditions and to deal with the equilibrium model, we assume that the market is in perfect competition, where players compete in terms of deciding their optimal quantities. Thus, $q_{jt\omega}$ also becomes a decision variable for the retailer. This implies that price tariff $P_{t}$ is considered exogenous for both retailers’ and consumers’ problems, and its equilibrium value will naturally come out as the market clearing price, i.e., the price that satisfies all the players’ (retailers’ and consumers’) optimality conditions.
The KKT optimality conditions derived from the retailer’s problem are:
@@
3.5 MILP formulation
We solve the equilibrium problem as MILP by concatenating the KKT optimality conditions (((3.3a))-((3.3h)) and ((3.5a))-((3.5k))). The complementarity slackness conditions ((3.3d))-((3.3f)) and ((3.5f))-((3.5i)) are linearized by introducing an auxiliary variable for each condition.
Then to exploit the efficient performance of MILP solvers, the resulting system of MIL constraints can be reformulated as a MILP by including an arbitrary constant as the objective function:

### v6

@@ 3 Problem formulation
Most electricity markets exhibit a hierarchical decision-making structure that allows participants with market power to potentially anticipate the reactions of subsequent decision-making entities. In this regard, bilevel programming is an adequate optimization-based modeling technique for representing such interactions. It can be defined in various forms, including single-leader single follower, single leader multi follower, multi leader single-follower, and multi leader multi follower games. While the single leader- single follower case has been extensively analyzed since von Stackelberg’s seminal work (Von Stackelberg, 1952), its extensions, accounting for multiple followers, remain less explored. This work falls into the category where we model the problem faced by a single retailer with market power acting as a leader and multiple electricity consumers as followers. This scenario is realistic from the perspective of a single retailer aiming to set adequate price tariffs for consumers. Similarly, since consumers are attached to a single retailer for their short-term decisions, we neglect potential interactions with rival retailers in capturing switching consumers. There are two common assumptions used to approach bilevel problems in the literature. i) The optimistic approach assumes that if a follower has more than one optimal solution, it will choose the solution that maximizes the leader’s objective. ii) For any fixed decision of the retailer on the tariff, each consumer optimizes their behavior independently, without considering the challenges faced by rival consumers (Kovács, 2019).
@@
$B_{jt\omega}$ quadratic coefficient of utility function for consumer $j$, hour $t$, scenario $\omega$ (€/kWh2)
@@
$\lambda_{j\omega}$ total quantity variation of demand for consumer $j$ under scenario $\omega$
@@
Thus, we assume that a single retailer wants to optimize its expected profit, and $\left|J\right|$ consumers want to minimize the difference between the cost of purchasing electricity and the utility of their consumption. That is, the retailer determines the selling prices to maximize its expected profit, which depends on the consumers’ purchases, and the consumers determine their load demands, which depend on the retail price tariffs. Consumers want to minimize disutility costs. First, we develop a bilevel programming problem assuming that retailers behave as leaders (upper-level) and that consumers behave as followers (lower-level). Then, we analyze the competitive equilibrium between the retailer and the consumers, assuming no market power.
@@
Bilevel programming with a strategic retailer is reformulated as an MPEC problem. This is done by recasting the consumers’ problems using their KKT optimality conditions, which are introduced as constraints in the retailer’s problem. Uncertainties related to consumer demand, spot market quantities and prices are addressed using stochastic programming.
@@
For the equilibrium model with perfect competition, we formulate an MILP problem by linearizing the complementarity conditions of the corresponding retailer and consumers’ KKT optimality conditions. In addition, an NLP reformulation of the equilibrium problem is derived using the methods proposed by Leyffer and Munson (2010), which provides a solution equivalent to the MILP formulation.
The model has first-stage and second-stage decisions. Prior to day-ahead market clearance, the retailer decides the price tariff for hour $t$ (first stage). However, the spot market price, the quantity purchased by the retailer from the electricity market, the quantity bought by consumers, and the electricity quantity imbalance decisions are second-stage and scenario-dependent variables. As indicated, we seek to model rational behavior for consumers, via a utility function whose marginal value captures a monotonic decrease in consumption with price, as is observed in many markets, including electricity. For its analytical simplicity, concavity, and interpretability, we chose a quadratic utility function of the form $U_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})=A_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})-$ $\frac{1}{2}B_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})^{2}$. i) It is simple to demonstrate market efficiency, which helps understand how well the allocation rule is designed to maximize the total utility of consuming $(q_{jt\omega}+\Delta_{jt\omega}^{C})$ at hour $t$ by consumer $j$ under scenario $\omega$. This is achieved by solving the optimal amount of electricity, where the marginal utility of consuming an additional unit is zero, i.e, $q_{jt\omega}+\Delta_{jt\omega}^{C}=\frac{A_{jt\omega}}{B_{jt\omega}}$. To this end, the efficient allocation of flexible and nonflexible electricity consumption for consumer $j$ at hour $t$ under scenario $\omega$ can be determined as $\Delta_{jt\omega}^{C}=\frac{A_{jt\omega}}{B_{jt\omega}}-q_{jt\omega}$ and $q_{jt\omega}=\frac{A_{jt\omega}}{B_{jt\omega}}-\Delta_{jt\omega}^{C}$, respectively. ii) Since $B_{jt\omega}>0$, $U^{{}^{\prime\prime}}_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})=-B_{jt\omega}$ is always negative, implying that the utility function is concave and demonstrates diminishing returns. This means that each additional unit adds less utility to consumer $j$ at hour $t$ under scenario $\omega$ than the previous unit. Furthermore, to demonstrate that the overall implications of our proposed models are not solely dependent on our modeling choice, we also compare the results using a linear utility function in our analysis.
@@ 3.1 Retailer’s problem
The retailer maximizes its expected profit subject to technical and economic constraints, and accounts for uncertainty, which is modeled by a discrete number of scenarios. The retailer’s problem is stated mathematically as follows:
@@
where $\bf{X}$=$\{P_{t},q_{t\omega}^{S},\delta_{t\omega},y_{t\omega}\}$ is the retailer’s set of decision variables, while $\mu_{t\omega},\alpha_{t\omega}^{+},\alpha_{t\omega}^{-}$, $\theta_{t\omega},\gamma_{t},\beta_{t\omega}$ are the dual variables associated with constraints ((3.1b))-((3.1g)).
The first term in the objective function represents the retailer’s revenue from the quantity bought by consumer $j$ at hour $t$ under scenario $\omega$. This quantity is multiplied by the price $P_{t}$, which is the retail price of electricity (tariff) and a first-stage decision set by the retailer for each hour $t$. The second term is the cost of purchasing the electricity quantity $q_{t\omega}^{S}$ at the spot market price $P_{t\omega}^{S}$ at hour $t$ and with scenario $\omega$. The last term represents a cost, penalizing electricity imbalances. This cost can be positive or negative, where the variable $y_{t\omega}$ denotes the absolute value of these imbalances. The parameter $\sigma_{\omega}$ represents the probability associated with scenario $\omega$. The summation of the product over index $\omega\in\Omega$ represents the expected profit, which is what the retailer wants to maximize over the time horizon.
Constraint ((3.1b)) is the electricity balancing constraint that guarantees the total demand by all consumers at hour $t$ with scenario $\omega$ minus the retailer’s supply, which is the imbalance at that particular time. The linear sets of constraints ((3.1c)), ((3.1d)) and ((3.1e)) ensure that $y_{t\omega}=|\delta_{t\omega}|$. These imbalances are penalized in the objective function with weight $C$. Constraints ((3.1f)) and ((3.1g)) are nonnegativity constraints on the spot market quantity and price decided by the retailer for tariffs at hour $t$, respectively. $P_{t}$ is a first-stage decision that needs to be settled before uncertainty realization, while $q_{jt\omega},q_{t\omega}^{S},\delta_{t\omega}$, $P_{t\omega}^{S}$ and $y_{t\omega}$ are second stage decisions. The dual variables for each group of constraints are indicated on the right side of the colons.
@@ 3.2 Consumers’ problem
In modeling consumer problems with DR, it is common to explicitly specify the type of consumer (residential, industrial, or commercial). This allows part of their consumption to be considered flexible, while another portion is inelastic and must be met at all times.
For instance, Halvgaard et al. (2012) model heat pumps for heating residential buildings, where the heating system of the house becomes flexible consumption in the smart grid. Zugno et al. (2013) extend this model by considering the heating dynamics of a building, the indoor temperature, or the temperature inside a water tank where flexibility (the output of interest) is characterized by the indoor temperature. They also consider the inflexible part of consumption in their model. However, in our model, including individual appliances would require a large number of additional variables and constraints. This would further complicate the numerical resolution and the economic interpretation of the results we aim to deliver. Moreover, we believe that the main demand patterns/profiles can be accounted for by the appropriate intertemporal adjustment of the utility function parameters.
By considering the consumers’ utility function explicitly in the objective function, we explore a residential consumer model with flexibility towards the dynamic price reported by the retailer. Thus, consumers face an economic problem where balancing the trade-off between electricity procurement costs and consumption utility is crucial. They must weigh the discomfort of deviating from their preferred consumption against the benefits of lower prices in a flexible pricing scheme. In particular, the consumers’ problem can be represented within a game-theoretical setting, as it is parametrized by the decision of the retailer (hourly price $P_{t}$). The problem for consumer $j$ with scenario $\omega$ can be expressed as:
@@
where the objective function is the minus social welfare (cost minus utility) and $q_{jt\omega}$ (inelastic demand) and $\Delta_{jt\omega}^{C}$ (flexible demand) are consumer’s decision variables. In the objective function $P_{t}q_{jt\omega}$ is the cost of purchasing $q_{jt\omega}$ electricity and $A_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})-\frac{1}{2}B_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})^{2}$ is the quadratic utility function that measures the benefit that consumer $j$ achieves by consuming the amount of energy ($q_{jt\omega}+\Delta_{jt\omega}^{C}$) during hour $t$, scenario $\omega$.
Considering the consumers’ objective function without any further constraints, consumption would only occur in periods when the real-time price is lower than the marginal benefit. Note that the marginal utility is a demand function $P_{t}=A_{jt\omega}-B_{jt\omega}(q_{jt\omega}+\Delta_{jt\omega}^{C})$ where the electricity quantity demanded decreases as the electricity price increases. In other words, at the optimum, the marginal utility (benefit) from consuming electricity equals the retail price tariff $P_{t}$.
Parameters $A_{jt\omega}$ and $B_{jt\omega}$ are the intercept and slope of the demand function, which are important parameters for characterizing the behavior of consumers toward the market outcomes of the model. Different values for these parameters can capture the dynamics of consumer demand. Constraint ((3.2c)) sets the lower and upper limits for flexible consumption, where $\Delta_{j}^{C_{max}}$ is the level of flexibility of consumer $j$. The inequalities guarantee that the flexible part of consumption at hour $t$ falls in the range between $-\Delta_{j}^{C_{max}}$ and $+\Delta_{j}^{C_{max}}$, which bound consumption increase and decrease, respectively. If $\Delta_{j}^{C_{max}}=0$ indicates that there is no flexibility in demand, while $\Delta_{j}^{C_{max}}>0$ indicates flexible demand. In the latter case, the consumer considers a dynamic price tariff that ensures cost savings and better welfare. In addition, constraint ((3.2d)) ensures that the net total flexible demand during the planning time horizon is zero. In the consumers’ problem, $q_{jt\omega}$ and $\Delta_{jt\omega}^{C}$ are second-stage decisions. Note that since the dynamic electricity price $P_{t}$ enters the consumers’ problem as a parameter (it is only a variable in the retailer problem), the optimization problems of consumers are convex (quadratic objective function with linear constraints). Therefore, we can replace each consumer’s problem with its corresponding KKT optimality conditions (Haghighat and Kennedy, 2012), which are sufficient for optimality.
@@ 3.2.1 KKT formulation of the consumer problem
where $\mathcal{L}_{j\omega}$ is the Lagrangian function for consumer $j$ and scenario $\omega$ of problem ((3.2)). By convention, the symbol $\perp$ indicates complementarity so that any of the two inequalities are satisfied as equality, i.e., the product of each expression and the corresponding dual variable must be zero. Note that the system of KKT optimality conditions is linear, with the exception of the complementarity conditions ((3.3d))-((3.3f)).
@@ 3.3 The Retailer’s MPEC problem
Figure 2: Game-theoretical framework for modeling demand response as a single leader-follower problem. The solid arrows indicate information signals from the retailer, while the dashed arrows represent inferences about consumers’ behavior.
@@
The second type of nonlinearities can be linearized using the strategy proposed in Fortuny-Amat and McCarl (1981), which involves applying the big-M technique. However, we still face the nonlinear product $P_{t}q_{jt\omega}$, which, as opposed to other problem settings Ruiz and Conejo (2009) and Sadat and Fan (2017), cannot be linearized by employing the strong duality theorem. However, due to the observed good numerical performance, we have opted to solve the original MPEC problem with a state-of-the-art NLP solver such as KNITRO (Artelys KNITRO, 2023).
@@ 3.4 The equilibrium model
In the case of perfect competition, we tackle the equilibrium model by jointly solving the KKT optimality conditions for retailers and consumers. This is achieved using two techniques: i) by linearizing and casting the KKT systems of equations as a MILP problem, employing the Big-M technique from Fortuny-Amat and McCarl (1981), and ii) by directly addressing and reformulating the problem as an NLP, following the approach in Leyffer and Munson (2010). However, we mainly present the reformulation of the equilibrium problem as an MILP and its results, for our analysis. For details on the reformulation of NLP problems, the interested reader is referred to Abate et al. (2021).
@@ 3.4.1 KKT formulation of the retailer problem
To derive the retailer’s KKT optimality conditions and address the equilibrium model, we assume a perfectly competitive market where players compete by choosing their optimal quantities. Consequently, $q_{jt\omega}$ also becomes a decision variable for the retailer. This setup implies that the price tariff $P_{t}$ is considered exogenous for both retailers and consumers. Its equilibrium value will naturally emerge as the market-clearing price, i.e., the price that satisfies the optimality conditions of all players (retailers and consumers).
The KKT optimality conditions derived from the retailer’s problem are as follows:
@@
3.5 Equilibrium problem formulation
Equilibrium problems in competitive markets can be approached in several ways. i) As an MILP by linearizing complementarity constraints with the Big-M method. ii) As an NLP, where products of the left- and right-hand sides of complementarity constraints form the objective function, with remaining feasibility and stationarity conditions as constraints. iii) By linearizing complementarity constraints using a special order set type 1 (SOS type 1) method, which avoids arbitrary constants but increases computational cost. iv) By formulating a mixed complementarity problem (MCP), which avoids Big-M values or nonlinear formulations by directly incorporating complementarity conditions and can be solved using tools like PATH or KNITRO. In this work, we choose the MILP approach with the Big-M method over MCP due to the broad availability and performance of MILP solvers, which align with practical industry applications. However, the Big-M method has limitations, as inappropriate values can introduce inaccuracies or computational issues (Pineda and Morales, 2019). To address this, we select Big-M values based on capacity constraints and use a trial-and-error approach for Big-M values associated with bounds on dual variables. We also implement an NLP reformulation to validate our MILP results, which are consistent across both methods. We solve the equilibrium problem as MILP by concatenating the KKT optimality conditions (((3.3a))-((3.3h)) and ((3.5a))-((3.5k))). The complementarity slackness conditions ((3.3d))-((3.3f)) and ((3.5f))-((3.5i)) are linearized by introducing an auxiliary variable for each condition.
Then, to exploit the efficient performance of MILP solvers, the resulting system of MIL constraints can be reformulated as an MILP by including an arbitrary constant as the objective function: